OpenStudy (anonymous):

Which of the following is the vertex of the function? y=x^2-6x+9 (3, 0) (-3, 36) (4, 1) (1, 4)

OpenStudy (anonymous):

To find the vertex we can use -b/2a where a = 1 b = -6 c = 9 What do you get for -b/2a?

OpenStudy (anonymous):

- - 6/ 2(1) = 6/2 = 3 Which means the answer is ...

OpenStudy (anonymous):

a .

OpenStudy (anonymous):

Excellent!